# Implicit Gaussian process representation of vector fields over arbitrary latent manifolds

### v1

@@ 1 Introduction
However, many datasets contain a richer structure comprising a smoothly varying vector field over the manifold. Prime examples are dissipative dynamical systems where post an initial transient phase, trajectories converge to a manifold in state space (Fefferman et al., 2016). Likewise, in neuroscience, smooth vector fields arise from the firing rate trajectories of neural populations evolving over neural manifolds, which is instrumental in neural information coding (Sussillo & Barak, 2013; Khona & Fiete, 2022; Gardner et al., 2022). Smooth vector fields are also pertinent in areas like gene expression profiling during development (La Manno et al., 2018) and multireference rotational alignment in cryoelectron microscopy (Singer & Wu, 2012). The breadth of these applications emphasises the need to generalise current learning paradigms to capture both the manifold structure and its associated vector field.
@@
To address this need, a promising avenue is to consider the Laplace-Beltrami operator as a hierarchy of Laplacians that act on tensor bundles of a manifold with increasing order. The first member of this hierarchy is the Laplace-Beltrami operator, which acts on rank-00 tensors, i.e., scalar signals. Similarly, higher-order signals, including vector fields, have associated Laplacian operators, which can encode their spatial regularity. Among these, the connection Laplacian (Barbero et al., 2022), defined on vector bundles, and the related sheaf Laplacian (Knöppel et al., 2013; Bodnar et al., 2022), which allows the vector spaces on nodes to have different dimensions, are emerging as leading operators in machine learning (Bronstein et al., 2017; Gosztolai et al., 2023). Both of these operators are related to heat diffusion of higher-order signals over manifolds (Singer & Wu, 2012; Sharp et al., 2019) and thus intrinsically encode the signals’ smoothness. The connection Laplacian is particularly appealing because it can be constructed, even when the manifold is unknown, from graph-based data descriptions (Singer & Wu, 2012; Budninskiy et al., 2019). We, therefore, asked how one could use this discrete approximation to derive continuous functions that implicitly represent the vector field over the manifold. Such representation could take account of the global regularity of the vector field to reconstruct intricate vector field structures lost in data sampling.
@@
Gaussian processes – a renowned family of nonparametric stochastic processes – offer an excellent framework for learning implicit functional descriptions of data. While GPs are traditionally defined on Euclidean spaces (Rasmussen & Williams, 2006), several studies have extended them to Riemannian manifolds. However, these studies have either considered scalar signals (Wilson et al., 2021; Mallasto & Feragen, 2018; Mallasto et al., 2020; Borovitskiy et al., 2020; 2021; Jensen et al., 2020) or vector signals Hutchinson et al. (2021) but only in cases where the underlying manifold is known and is analytically tractable, such as spheres and tori. In this work, we generalise GPs to vector fields on arbitrary latent manifolds, which can only be approximated based on local similarities between data points, making them applicable in practice to real-world datasets.
@@
Our contributions are as follows. (i) We generalise GPs to vector-valued data using the connection Laplacian operator, assuming that the data originates from a stationary stochastic process. (ii) We show that the resulting Riemannian manifold vector field GP (RVGP) method encodes the manifold and vector field’s smoothness as inductive biases, enabling out-of-sample predictions from sparse or obscured data. (iii) To underscore the practical implications of RVGP, we apply it to electroencephalography (EEG) recordings from both healthy individuals and Alzheimer’s disease patients. The global spatial regularity learnt by our method significantly outperforms the state-of-the-art approaches for reconstructing high-density electrical fields from low-density EEG arrays. This enables our method to better resolve vector field singularities and dramatically increase the classification power of disease states.
Taken together, our technique opens the door to a differential geometric formulation of kernel-based operators and demonstrates a direct relevance for fundamental and clinical neuroscience.
@@ 2 Comparison with related works
Several closely related works in the GP literature have provided various definitions of GPs on Riemannian manifolds. One line of works defined GPs as manifold-valued processes $f:{\mathbb{X}}\to\mathcal{M}$ (Mallasto & Feragen, 2018; Mallasto et al., 2020) by using the exponential map of the manifold to perform regression in the tangent space. However, these works require that the manifold $\mathcal{M}$ be known in order to define the exponential map. More notable are studies which define GPs as scalar-valued functions $f:{\mathbb{X}}\to\mathbb{R}$. For example, considering data points as samples from a stationary stochastic process, the domain ${\mathbb{X}}$ can be defined based on positional encoding using eigenfunctions of the Laplace-Beltrami operator (Solin & Särkkä, 2020; Borovitskiy et al., 2020) or the graph Laplacian (Borovitskiy et al., 2021). However, these works cannot be directly applied to vector-valued signals by treating vector entries as scalar channels. This is because these channels are generally not independent but related through the curvature of the manifold. To address this gap, Hutchinson et al. (2021) defined GPs as functions $f:{\mathbb{X}}\to\mathcal{TM}$ over the tangent bundle $\mathcal{TM}$ by first mapping the manifold isometrically to Euclidean space and then using a multi-input, multi-output GP to learn the projected signal. However, Hutchinson et al. (2021) has focused on cases when the manifold is explicitly known, and its mapping to Euclidean space can be explicitly defined. Here, we are specifically interested in the case where the manifold is unknown – a common scenario in many scientific domains. To achieve this, we generalise the works of Borovitskiy et al. (2020; 2021) to vector fields on Riemannian manifolds using the connection Laplacian operator and its eigenvectors as positional encoding.
@@ 3.1 Gaussian processes in Euclidean spaces
A GP is a stochastic process $f:{\mathbb{X}}\to\mathbb{R}^{d}$ defined over a set ${\mathbb{X}}$ such that for any finite set of samples $X=({\bm{x}}_{1},\dots,{\bm{x}}_{n})\in{\mathbb{X}}^{n}$, the random variables $(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{n}))\in\mathbb{R}^{n\times d}$ are jointly multivariate Gaussian, $\mathcal{N}({\bm{x}};{\bm{\mu}},{\bm{K}})$, with mean vector ${\bm{\mu}}=m(X)$ and covariance matrix ${\bm{K}}=k(X,X)$. Consequentially, a GP is fully characterised by its mean function $m({\bm{x}}):=\mathbb{E}(f({\bm{x}}))$ and covariance function $k({\bm{x}},{\bm{x}}^{\prime}):=\mathrm{Cov}(f({\bm{x}}),f({\bm{x}}^{\prime}))$, also known as kernel, and denoted as $f\sim\mathcal{GP}(m({\bm{x}}),k({\bm{x}},{\bm{x}}^{\prime}))$. It is typical to assume that ${\bm{m}}({\bm{x}})=0$, which does not reduce the expressive power of GPs (Rasmussen & Williams, 2006).
@@ 4 Intrinsic representation of vector fields over arbitrary latent manifolds
Figure 1: Construction of vector-valued Gaussian processes on unknown manifolds. A Input data consists of samples from a vector field over a latent manifold. B The manifold is approximated by a proximity graph. Black circles mark two sample points, $i$ and $j$ and their graph neighbourhood. C The tangent bundle is a collection of locally Euclidean vector spaces over the manifold. It is approximated by parallel transport maps between local tangent space approximations. D The eigenvectors of the connection Laplacian are used as positional encoding to define the GP that leans the vector fields. E The GP is evaluated as unseen points to predict the smoothest possible vector field that is consistent with the training data. We use this GP to accurately predict singularities, near which sampling is typically sparse.
@@ 4.1 Vector-valued GPs on unknown manifolds
We consider training data consisting of pairs $\{({\bm{x}}_{i},{\bm{v}}_{i})|i=1,\dots,n\}=(X,V)$, where ${\bm{x}}_{i}$ are samples from a manifold $\mathcal{M}\subset\mathbb{R}^{d}$ and ${\bm{v}}_{i}$ are sampled from the tangent bundle $\mathcal{TM}=\cup_{i}\mathcal{T}_{i}\mathcal{M}$. If the dimension of the manifold is $m\leq d$ the tangent spaces $\mathcal{T}_{i}\mathcal{M}:=\mathcal{T}_{{\bm{x}}_{i}}\mathcal{M}$ anchored to ${\bm{x}}_{i}$ are isomorphic as a vector space to $\mathbb{R}^{m}$. Importantly, we assume that both $\mathcal{M}$ and $\mathcal{T}_{i}\mathcal{M}$ are unknown and seek a GP to provide an implicit description of the vector field over $\mathcal{M}$ that agrees with the training set and provides a continuous interpolation at out-of-sample test points with controllable smoothness properties.
@@
We first fit a proximity graph $G=(V,E)$ to $X$, defined based on some notion of similarity (spatial or otherwise) in the data. While $G$ approximates $\mathcal{M}$ it will not restrict the domain to $V$ as in Borovitskiy et al. (2021). Then, to approximate $\mathcal{TM}$, note that the tangent spaces do not come with preferred coordinates. However, being isomorphic to $\mathbb{R}^{m}$, $\mathcal{T}_{i}\mathcal{M}$ can be parametrised by $m$ orthogonal vectors in the ambient space $\mathbb{R}^{d}$, to form a local frame, or gauge, $\mathbb{T}_{i}$. To obtain this frame, we take vectors $\mathbf{e}_{ij}\in\mathbb{R}^{d}$ from $i$ to $N$ nearest nodes $j$, assuming that they span $\mathcal{T}_{i}\mathcal{M}$ [^2] In practice, we pick $N$ closest nodes to $i$ on the proximity graph where $N$ is a hyperparameter. Larger $N$ increases the overlaps between the nearby tangent spaces. We find that $N=2D_{ii}$ is often a good compromise between locality and robustness to noise of the tangent space approximation. and form a matrix by stacking them column-wise. The left singular vectors corresponding to the $m$ largest singular values yield the desired frame
@@
Then, $\hat{{\bm{v}}}_{i}=\mathbb{T}^{T}_{i}{\bm{v}}_{i}$ acts as a projection of the signal to the tangent space in the $\ell_{2}$ sense.
@@
The notion of vector field smoothness is formalised by the parallel transport map $\mathcal{P}_{j\to i}$ that aligns $\mathcal{T}_{j}\mathcal{M}$ with $\mathcal{T}_{i}\mathcal{M}$ to allow the comparison of vectors in a common space. While parallel transport is generally path dependent, we assume that $i,j$ are close enough such that $\mathcal{P}_{j\to i}$ is the unique smallest rotation. Indeed, the construction of the nearest neighbour proximity graph limits pairs $i,j$ to be close in space. This is known as the Lévy-Civita connection and can be computed as a matrix $\mathbf{O}_{ji}\in O(m)$ in the orthogonal group (rotation and reflection)
@@
where $||\cdot||_{F}$ is the Frobenius norm and can be uniquely computed by the Kabsch algorithm in $\mathcal{O}(m)$-time (Kabsch, 1976).
Using the parallel transport maps, we may define the connection Laplacian (Singer & Wu, 2012), a block matrix ${\bm{L}}_{c}\in\mathbb{R}^{nd\times nd}$, whose $(i,j)$ block entry is given by
@@
Let us remark that ${\bm{L}}_{c}$ prescribes the smoothness of the vector field over an underlying continuous manifold that agrees with the available training data. In fact, as $n\to\infty$, the eigenvectors of ${\bm{L}}_{c}$ converge to the eigenfunctions of the connection Laplacian over the tangent bundle (Singer & Wu, 2017), such that the corresponding continuous signal satisfies the associated vector diffusion process (Berline N., 1996). The solution of this diffusion process, minimises the vector Dirichlet energy $\sum_{ij\in E}w_{ij}|{\bm{v}}_{i}-{\bm{O}}_{ij}{\bm{v}}_{j}|^{2}$, which quantifies the smoothness of the vector field (Knöppel et al., 2015).
@@
We are now ready to define a GP to regress the vector field over $\mathcal{M}$. To this end, we consider a positional encoding of points on the tangent bundle $\mathcal{TM}$ based on the spectrum of the connection Laplacian, ${\bm{L}}_{c}={\bm{U}}_{c}{\bm{\Lambda}}_{c}{\bm{U}}_{c}^{T}$, where ${\bm{\Lambda}}_{c},{\bm{U}}_{c}\in\mathbb{R}^{nm\times nm}$. Comparing with Eq. 7 one notices that each node represents a vector space of dimension $m$. Using the eigenvectors corresponding to the $k$ smallest eigenvalues, and for some point ${\bm{x}}$ corresponding to node $i$, we consider the matrix
@@
$({\bm{U}}_{c})_{i}=\sqrt{nm}\begin{pmatrix}u_{im,1}&\dots&u_{im,k}\\ \vdots&&\vdots\\ u_{(i+1)m,1}&\dots&u_{(i+1)m,k}\end{pmatrix}\in\mathbb{R}^{m\times k},$ (13)
@@
obtained by keeping $k$ columns of ${\bm{U}}_{c}$. This allows us to define the positional encoding of ${\bm{x}}$ by mapping the eigencoordinates defined in the respective tangent spaces using ${\mathbb{T}}_{i}$ back into the ambient space
${\bm{P}}_{{\bm{x}}}={\mathbb{T}}_{i}({\bm{U}}_{c})_{i}\in\mathbb{R}^{d\times k}.$ (14)
@@
Then, by analogy to Eqs. 5-6, we define the scalar-valued Matérn GP that represents the vector field over the manifold as $f_{\mathcal{TM}}:\mathcal{TM}\to\mathbb{R}^{d}$ with a $\mathbb{R}^{d\times d}$-valued kernel
$k_{\mathcal{TM}}({\bm{v}},{\bm{v}}^{\prime})=\sigma^{2}{\bm{P}}_{{\bm{v}}}\Phi({\bm{\Lambda}}_{c})^{-2}{\bm{P}}_{{\bm{v}}^{\prime}}^{T},$ (15)
@@
where ${\bm{v}},{\bm{v}}^{\prime}\in\mathcal{TM}$. Note that we recover $f_{\mathcal{M}}$ for scalar signals ($m=1$), where Laplace-Beltrami operator equals the connection Laplacian for a trivial line bundle on $\mathcal{M}$. Therefore, the tangent spaces become trivial (scalar), recovering the well-known Laplacian eigenmaps ${\bm{P}}_{{\bm{x}}}=\sqrt{n}(u_{i,1},\dots,u_{i,k})$. Likewise, $k_{\mathcal{TM}}({\bm{v}},{\bm{v}}^{\prime})$ reduces to the scalar-valued kernel $k_{\mathcal{M}}({\bm{x}},{\bm{x}}^{\prime})$ for ${\bm{x}},{\bm{x}}^{\prime}\in\mathcal{M}$ in Eq. 6. Note that the kernel in Eq. 15 is a member of a general class defined in Hutchinson et al. (2021, Proposition 7) arising as an isometric projection of the tangent bundle into Euclidean space. Our construction makes this projection explicit using the connection Laplacian positional encoding (Eq. 13) and tangent spaces (Eq. 10). In Fig. 2, we show that this construction leads to specific performance advantages in recovering vector field singularities.

### v2

@@ 1 Introduction
However, many datasets contain a richer structure comprising a smoothly varying vector field over the manifold. Prime examples are dissipative dynamical systems where, post an initial transient phase, trajectories converge to a manifold in state space (Fefferman et al., 2016). Likewise, in neuroscience, smooth vector fields arise from the firing rate trajectories of neural populations evolving over neural manifolds, which is instrumental in neural information coding (Sussillo & Barak, 2013; Khona & Fiete, 2022; Gardner et al., 2022). Smooth vector fields are also pertinent in areas like gene expression profiling during development (La Manno et al., 2018) and multireference rotational alignment in cryoelectron microscopy (Singer & Wu, 2012). These applications emphasise the need to generalise current learning paradigms to capture both the manifold structure and its associated vector field.
@@
To address this need, a promising avenue is to consider the Laplace-Beltrami operator as a hierarchy of Laplacians that act on tensor bundles of a manifold with increasing order. The first member of this hierarchy is the Laplace-Beltrami operator, which acts on rank-00 tensors, i.e., scalar signals. Similarly, higher-order signals, including vector fields, have associated Laplacian operators, which can encode their spatial regularity. Among these, the connection Laplacian (Barbero et al., 2022), defined on vector bundles, and the related sheaf Laplacian (Knöppel et al., 2013; Bodnar et al., 2022), which allows the vector spaces on nodes to have different dimensions, are emerging as leading operators in machine learning (Bronstein et al., 2017; Battiloro et al., 2023; Gosztolai et al., 2023). These operators are related to heat diffusion of higher-order signals over manifolds (Singer & Wu, 2012; Sharp et al., 2019) and thus intrinsically encode the signals’ smoothness. The connection Laplacian is particularly appealing because it can be constructed, even when the manifold is unknown, from graph-based data descriptions (Singer & Wu, 2012; Budninskiy et al., 2019). We, therefore, asked how one could use this discrete approximation to derive continuous functions that implicitly represent the vector field over the manifold. Such representation could use the global regularity of the vector field to reconstruct intricate vector field structures lost in data sampling.
@@
Gaussian processes – a renowned family of nonparametric stochastic processes – offer an excellent framework for learning implicit functional descriptions of data. While GPs are traditionally defined on Euclidean spaces (Rasmussen & Williams, 2006), several studies have extended them to Riemannian manifolds. However, these studies have either considered scalar signals (Wilson et al., 2021; Mallasto & Feragen, 2018; Mallasto et al., 2020; Borovitskiy et al., 2020; 2021; Jensen et al., 2020) or vector signals Hutchinson et al. (2021) but only in cases where the underlying manifold is known and is analytically tractable, such as spheres and tori. In this work, we generalise GPs to vector fields on arbitrary latent manifolds, which can only be approximated based on local similarities between data points, making them applicable to real-world datasets.
@@
Our contributions are as follows. (i) We generalise GPs to vector-valued data using the connection Laplacian operator, assuming that the data originates from a stationary stochastic process. (ii) We show that the resulting Riemannian manifold vector field GP (RVGP) method encodes the manifold and vector field’s smoothness as inductive biases, enabling out-of-sample predictions from sparse or obscured data. (iii) To underscore the practical implications of RVGP, we apply it to electroencephalography (EEG) recordings from both healthy individuals and Alzheimer’s disease patients. The global spatial regularity learnt by our method significantly outperforms the state-of-the-art approaches for reconstructing high-density electrical fields from low-density EEG arrays. This enables our method to better resolve vector field singularities and dramatically increase the classification power of disease states. In sum, our work enables a differential geometric formulation of kernel-based operators and demonstrates a direct relevance for fundamental and clinical neuroscience.
@@ 2 Comparison with related works
Several closely related works in the GP literature have provided various definitions of GPs on Riemannian manifolds. One line of works defined GPs as manifold-valued processes $f:{\mathbb{X}}\to\mathcal{M}$ (Mallasto & Feragen, 2018; Mallasto et al., 2020) by using the exponential map of the manifold to perform regression in the tangent space. However, these works require that the manifold $\mathcal{M}$ be known to define the exponential map. More notable are studies which define GPs as scalar-valued functions $f:{\mathbb{X}}\to\mathbb{R}$. For example, considering data points as samples from a stationary stochastic process, the domain ${\mathbb{X}}$ can be defined based on positional encoding using eigenfunctions of the Laplace-Beltrami operator (Solin & Särkkä, 2020; Borovitskiy et al., 2020) or the graph Laplacian (Borovitskiy et al., 2021). However, these works cannot be directly applied to vector-valued signals by treating vector entries as scalar channels. This is because these channels are generally not independent but related through the curvature of the manifold. To address this gap, Hutchinson et al. (2021) defined GPs as functions $f:{\mathbb{X}}\to\mathcal{TM}$ over the tangent bundle $\mathcal{TM}$ by first mapping the manifold isometrically to Euclidean space and then using a multi-input, multi-output GP to learn the projected signal. However, Hutchinson et al. (2021) has focused on cases when the manifold is explicitly known, and its mapping to Euclidean space can be explicitly defined. Here, we are specifically interested in the case where the manifold is unknown – a common scenario in many scientific domains. To achieve this, we generalise the works of Borovitskiy et al. (2020; 2021) to vector fields on Riemannian manifolds using the connection Laplacian operator and its eigenvectors as positional encoding.
@@ 3.1 Gaussian processes in Euclidean spaces
A GP is a stochastic process $f:{\mathbb{X}}\to\mathbb{R}^{d}$ defined over a set ${\mathbb{X}}$ such that for any finite set of samples $X=({\bm{x}}_{1},\dots,{\bm{x}}_{n})\in{\mathbb{X}}^{n}$, the random variables $(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{n}))\in\mathbb{R}^{n\times d}$ are jointly multivariate Gaussian, $\mathcal{N}({\bm{x}};{\bm{\mu}},{\bm{K}})$, with mean vector ${\bm{\mu}}=m(X)$ and covariance matrix ${\bm{K}}=k(X,X)$. Consequentially, a GP is fully characterised by its mean function $m({\bm{x}}):=\mathbb{E}(f({\bm{x}}))$ and covariance function $k({\bm{x}},{\bm{x}}^{\prime}):=\mathrm{Cov}(f({\bm{x}}),f({\bm{x}}^{\prime}))$, also known as the kernel, and denoted as $f\sim\mathcal{GP}(m({\bm{x}}),k({\bm{x}},{\bm{x}}^{\prime}))$. It is typical to assume that ${\bm{m}}({\bm{x}})=0$, which does not reduce the expressive power of GPs (Rasmussen & Williams, 2006).
@@ 4 Intrinsic representation of vector fields over arbitrary latent manifolds
Figure 1: Construction of vector-valued Gaussian processes on unknown manifolds. A Input consists of samples from a vector field over a latent manifold $\mathcal{M}$. B The manifold is approximated by a proximity graph. Black circles mark two sample points, $i$ and $j$ and their graph neighbourhood. C The tangent bundle is a collection of locally Euclidean vector spaces over $\mathcal{M}$. It is approximated by parallel transport maps between local tangent space approximations. D The eigenvectors of the connection Laplacian are used as positional encoding to define the GP that learns the vector field. E The GP is evaluated as unseen points to predict the smoothest vector field that is consistent with the training data. We use this GP to accurately predict singularities, where sampling is typically sparse.
@@ 4.1 Vector-valued GPs on unknown manifolds
We consider training data consisting of pairs $\{({\bm{x}}_{i},{\bm{v}}_{i})|i=1,\dots,n\}$, where ${\bm{x}}_{i}$ are samples from a manifold $\mathcal{M}\subset\mathbb{R}^{d}$ and ${\bm{v}}_{i}$ are sampled from the tangent bundle $\mathcal{TM}=\cup_{i}\mathcal{T}_{i}\mathcal{M}$. If the dimension of the manifold is $m\leq d$ the tangent spaces $\mathcal{T}_{i}\mathcal{M}:=\mathcal{T}_{{\bm{x}}_{i}}\mathcal{M}$ anchored to ${\bm{x}}_{i}$ are isomorphic as a vector space to $\mathbb{R}^{m}$. Importantly, we assume that both $\mathcal{M}$ and $\mathcal{T}_{i}\mathcal{M}$ are unknown and seek a GP to provide an implicit description of the vector field over $\mathcal{M}$ that agrees with the training set and provides a continuous interpolation at out-of-sample test points with controllable smoothness properties.
@@
We first fit a proximity graph $G=(V,E)$ to $X$, defined based on some notion of similarity (spatial or otherwise) in the data. While $G$ approximates $\mathcal{M}$, it will not restrict the domain to $V$ as in Borovitskiy et al. (2021). Then, to approximate $\mathcal{TM}$, note that the tangent spaces do not have preferred coordinates. However, being isomorphic to $\mathbb{R}^{m}$, $\mathcal{T}_{i}\mathcal{M}$ can be parametrised by $m$ orthogonal vectors in the ambient space $\mathbb{R}^{d}$, to form a local frame, or gauge, $\mathbb{T}_{i}$. To obtain this frame, we take vectors $\mathbf{e}_{ij}\in\mathbb{R}^{d}$ from $i$ to $N$ nearest nodes $j$, assuming that they span $\mathcal{T}_{i}\mathcal{M}$ [^2] In practice, we pick $N$ closest nodes to $i$ on the proximity graph where $N$ is a hyperparameter. Larger $N$ increases the overlaps between the nearby tangent spaces. We find that $N=2D_{ii}$ is often a good compromise between locality and robustness to noise of the tangent space approximation. and form a matrix by stacking them column-wise. The left singular vectors corresponding to the $m$ largest singular values yield the desired frame
@@
Then, $\hat{{\bm{v}}}_{i}=\mathbb{T}^{T}_{i}{\bm{v}}_{i}$ acts as a projection of the signal to the tangent space in the $\ell_{2}$ sense. Note that $m$ does not need to be known ahead of time but is estimated as the average of the dominant single values across all estimated tangent spaces, e.g, based on a cutoff.
@@
The notion of vector field smoothness is formalised by the parallel transport map $\mathcal{P}_{j\to i}$ that aligns $\mathcal{T}_{j}\mathcal{M}$ with $\mathcal{T}_{i}\mathcal{M}$ to allow the comparison of vectors in a common space. While parallel transport is generally path dependent, we assume that $i,j$ are close enough such that $\mathcal{P}_{j\to i}$ is the unique smallest rotation. Indeed, constructing the nearest neighbour proximity graph limits pairs $i,j$ to be close in space. This is known as the Lévy-Civita connection and is computed as a matrix $\mathbf{O}_{ji}\in O(m)$ in the orthogonal group (rotation and reflection)
@@
where $||\cdot||_{F}$ is the Frobenius norm and is uniquely computable in $\mathcal{O}(m)$-time (Kabsch, 1976).
Using the parallel transport maps, we define the connection Laplacian (Singer & Wu, 2012), a block matrix ${\bm{L}}_{c}\in\mathbb{R}^{nm\times nm}$, whose $(i,j)$ block entry is given by
@@
Let us remark that ${\bm{L}}_{c}$ prescribes the smoothness of the vector field over an underlying continuous manifold that agrees with the available training data. In fact, as $n\to\infty$, the eigenvectors of ${\bm{L}}_{c}$ converge to the eigenfunctions of the connection Laplacian over the tangent bundle (Singer & Wu, 2017), such that the corresponding continuous signal satisfies the associated vector diffusion process (Berline N., 1996). The solution of this diffusion process minimises the vector Dirichlet energy $\sum_{ij\in E}w_{ij}|{\bm{v}}_{i}-{\bm{O}}_{ij}{\bm{v}}_{j}|^{2}$, which quantifies the smoothness of the vector field (Knöppel et al., 2015).
@@
We can now define a GP to regress the vector field over $\mathcal{M}$. To this end, we consider a positional encoding of points on the tangent bundle $\mathcal{TM}$ based on the spectrum of the connection Laplacian, ${\bm{L}}_{c}={\bm{U}}_{c}{\bm{\Lambda}}_{c}{\bm{U}}_{c}^{T}$, where ${\bm{\Lambda}}_{c},{\bm{U}}_{c}\in\mathbb{R}^{nm\times nm}$. Compared with Eq. 7, where each node corresponds to a point, each node now represents a vector space of dimension $m$. Thus, the positional encoding of some vector ${\bm{v}}$ at node $i$ is given by an $\mathbb{R}^{m\times k}$ matrix, rather than an $\mathbb{R}^{1\times k}$ vector, whose columns are the eigenvectors corresponding to the $k$ smallest eigenvalues in ${\bm{\Lambda}}_{c}$ and rows are the coordinates of $\mathcal{T}_{i}\mathcal{M}$:
@@
$(\widetilde{{\bm{U}}}_{c})_{i}=\sqrt{nm}\begin{pmatrix}u_{im,1}&\dots&u_{im,k}\\ \vdots&&\vdots\\ u_{(i+1)m,1}&\dots&u_{(i+1)m,k}\end{pmatrix}\in\mathbb{R}^{m\times k}.$ (13)
@@
The matrix $(\widetilde{{\bm{U}}}_{c})_{i}$ can also be thought of as the $i$-th slice of an $\mathbb{R}^{n\times m\times k}$ tensor. This allows us to define the positional encoding of ${\bm{v}}$ by mapping the eigencoordinates defined in the respective tangent spaces using ${\mathbb{T}}_{i}$ back into the ambient space
${\bm{P}}_{{\bm{v}}}={\mathbb{T}}_{i}(\widetilde{{\bm{U}}}_{c})_{i}\in\mathbb{R}^{d\times k}.$ (14)
@@
Then, by analogy to Eqs. 5-6, we define the vector-valued Matérn GP $f_{\mathcal{TM}}:\mathcal{TM}\to\mathbb{R}^{d}$ representing the vector field over the manifold with a $\mathbb{R}^{d\times d}$-valued kernel (Fig. 2A, top)
$k_{\mathcal{TM}}({\bm{v}},{\bm{v}}^{\prime})=\sigma^{2}{\bm{P}}_{{\bm{v}}}\Phi(\widetilde{{\bm{\Lambda}}}_{c})^{-2}{\bm{P}}_{{\bm{v}}^{\prime}}^{T},$ (15)
@@
where ${\bm{v}},{\bm{v}}^{\prime}\in\mathcal{TM}$ and $\widetilde{{\bm{\Lambda}}}_{c}=({\bm{\Lambda}}_{c})_{1:k,1:k}$. Note that we recover $f_{\mathcal{M}}$ for scalar signals ($m=1$), where Laplace-Beltrami operator equals the connection Laplacian for a trivial line bundle on $\mathcal{M}$. Therefore, the tangent spaces become trivial (scalar), recovering the well-known Laplacian eigenmaps ${\bm{P}}_{{\bm{x}}}=\sqrt{n}(u_{i,1},\dots,u_{i,k})$. Likewise, $k_{\mathcal{TM}}({\bm{v}},{\bm{v}}^{\prime})$ reduces to the scalar-valued kernel $k_{\mathcal{M}}({\bm{x}},{\bm{x}}^{\prime})$ for ${\bm{x}},{\bm{x}}^{\prime}\in\mathcal{M}$ in Eq. 6. However, $k_{\mathcal{M}}({\bm{x}},{\bm{x}}^{\prime})$ and $k_{\mathcal{TM}}({\bm{v}},{\bm{v}}^{\prime})$ are not linearly related because the underlying Laplace-Beltrami and connection Laplacian operators are linked by the curvature of the manifold as given by the Weitzenböck identity. Note that the kernel in Eq. 15 differs from that defined in Hutchinson et al. (2021), which considers an isometric projection of the tangent bundle into Euclidean space and constructs a scalar-valued kernel therein (Fig. 2A, bottom). Instead of approximating this isometric projection, which may be challenging for unknown manifolds, our constructive approach uses an intrinsic parametrisation of the manifold using only local similarities. This yields a matrix-valued kernel (Fig. 2A, top), which accounts for alignment of the vector field to principal curvature directions over $\mathcal{M}$. This is made explicit using the connection Laplacian positional encoding (Eq. 13) and tangent spaces (Eq. 10). In Sect. 5, we show that this construction leads to specific performance advantages in recovering vector field singularities.